# Lost in Translation, Found in Context: Sign Language Translation with Contextual Cues

### v1

@@ 2 Related Work
Continuous Sign Language Recognition (CSLR) typically involves recognising gloss sequences – spoken language words corresponding to semantic labels for individual signs – from a continuous signing video clip. It is crucial to distinguish CSLR from translation due to the significant grammatical differences between signed and spoken languages [67]. Benchmarks such as PHOENIX14 [39] and CSL-Daily [85], which have manually glossed training data, are widely used for CSLR. Due to lack of temporal alignment between signs and video frames, many studies [20, 18, 73, 35, 10, 87, 30, 58, 51, 83] utilise the Connectionist Temporal Classification (CTC) loss [27]. More recently, [60] demonstrates the advantages of using video-to-text retrieval on the challenging BOBSL dataset [2].
@@
Recent works also incorporate large-scale language foundation models. For example, [66] inputs pseudo-glosses directly into ChatGPT. [70, 62, 61, 36, 82] all fine-tune a pre-trained T5 model for SLT. [70] feeds in 3D landmark embeddings, while [62] and [61] use visual features from a BEVT-pre-trained [72] ISLR model and a MAE-pretained video encoder, respectively. Sign2GPT [75] leverages large-scale pre-trained vision (DINOv2 [55]) and language (XGLM [47]) models, incorporating adapters (LoRA [32]) for transfer to sign language. Similarly to these works, we make use of a pre-trained LLM (Llama3 [24]) and fine-tune it for SLT.
@@ 3.2 Model inputs
Visual features. Following [60, 57], we employ a Video-Swin model [48] pre-trained for ISLR (classification of isolated signs) to obtain our visual features. Specifically, we utilise the recent and strong model provided by [60], which is trained on the BOBSL videos [2] for a vocabulary of 8,697 signs using the automatic sign spottings from [52]. This ISLR model processes short video clips of 16 frames (i.e. less than 1 second in 25 fps videos) to produce a single 768-dimensional embedding vector. Specifically, this vector comes from spatio-temporal averaging of the penultimate layer output of Video-Swin. To capture fine-grained temporal details, we feed 16 consecutive frames into the sign video encoder and apply a sliding window with a stride of to obtain features , where represents the number of visual features per sentence. For example, when =2, we have on average 56 features (i.e. 4.5 seconds) in BOBSL sentences. Note that for experiments conducted on the How2Sign dataset [22], we further fine-tune the Video-Swin model using the spotting annotations provided by [23] for a vocabulary of 1,887 signs (see Sec. A.5 for details).
Pseudo-glosses. To represent sign sequences in text form, we apply the ISLR model mentioned above in a sliding window manner to record the classification predictions, and obtain pseudo-glosses , representing a sequence of words (or phrases) from the vocabulary of the classifier (e.g. 8k signs). Note that these sign category predictions are noisy, often including more labels than the number of signs occurring in the sentence video, and many false positives, which we wish to suppress via our LLM tuning (see Fig. 2). Our pseudo-glosses are similar to [60], except we only apply repetition grouping, but we do not filter out low-confidence annotations with a threshold – this allows the LLM to learn which ones are relevant. In contrast to [60] that uses pseudo-glosses for supervision, we simply employ them as additional inputs to our SLT model. Typically, there are around 22 glosses per BOBSL sentence (which is less than the number of visual features – 56 features on average).
@@ 3.3 Tuning the LLM with multifarious cues
Mapping network for visual features. All our inputs are already in text form except the visual features, which need a projection to map them into the text space, so that they can be fed into the pre-trained LLM. To this end, we train a simple mapping network, a 2-layer MLP with GELU [31] activation in between, projecting the visual features (of dimensionality 768) to the dimensionality of the LLM input embeddings (i.e. 4,096). We note that we add 1D temporal convolution layer to the mapping network when experimenting with the How2Sign dataset [22] (see Sec. A.5 for further details).
@@ 4 Experiments
In this section, we first present the datasets and a suite of evaluation protocols used in our experiments (Sec. 4.1), as well as baseline descriptions (Sec. 4.2). Next, we ablate various components of our framework (Sec. 4.3). We then show our improved translation performance compared to the state of the art on two challenging open-vocabulary benchmarks (Sec. 4.4). Finally, we illustrate qualitative results and discuss limitations (Sec. 4.5)
@@ 4.2 Baselines
Oracle: Sincan [65] (Vid+Prev). This oracle baseline is a multi-modal variant of Sincan [65], where the three previous ground truth sentences are also fed as additional context to the decoder at both training and inference times.
@@ 4.4 Comparison to the state of the art
BOBSL. We evaluate our model on the BOBSL test set, Sent-Test, using our full suite of evaluation metrics. We show in Tab. 3 that our approach gives a significant boost in all metrics compared to previous works. Specifically, compared to state-of-the-art approaches GFSLT [84] and Sign2GPT [75], we achieve a considerable increase in performance (40.3 vs 35.2 B-RT)—highlighting both the effectiveness of our method which leverages context and the increased challenge posed by the BOBSL dataset (as opposed to PHOENIX14T where [84, 75] were originally evaluated).
In the Oracle setup (the bottom block of Tab. 3), we compare to the setting of [65], where models have access to ground truth previous sentence and spottings extracted from the ground truth current sentence. When using the ground truth previous sentence at inference, our model outperforms [65] by a large margin (41.6 vs 35.8 B-RT). When replacing our pseudo-glosses with spottings, we further increase the margin, substantially outperforming their method (47.1 vs 37.0 B-RT, i.e. +10 point improvement).
@@ 5 Conclusion
In this work, we show that leveraging contextual information significantly enhances SLT performance in an open-vocabulary setting. Specifically, our framework utilises background descriptions from a captioning model and predictions of previous sentences, combined with pseudo-glosses and visual features. Through an extensive ablation study, we analyse the individual impact of each cue on sign language translation, and benchmark our method against previous state-of-the-art approaches on the BOBSL dataset to demonstrate its effectiveness. While our approach considerably outperforms prior methods, there remains a room for improvement in the translation quality to be deployed in real-world applications, necessitating to address the limitations mentioned above.
Acknowledgments. The images in this paper are used with kind permission of the BBC. This work was granted access to the HPC resources of IDRIS under the allocation 2024-AD011013395 made by GENCI. The authors would like to acknowledge the ANR project CorVis ANR-21-CE23-0003-01, a Google Research Scholar Award, and a Royal Society Research Professorship RPR1191132. The authors also thank Charles Raude, Prajwal KR, Prasanna Sridhar, Joon Son Chung, Makarand Tapaswi, Syrine Kalleli, and David Picard for their help.
@@ Appendix B Additional Experiments
We examine performance variations when using different LLM decoders (Sec. B.1), evaluate all possible cue combinations (Sec. B.2), and investigate scenarios with missing cues (Sec. B.3). We also showcase the performance of our ISLR backbone on the HowSign dataset (Sec. B.4), and report the reproduction results of GFSLT [84] and Sign2GPT [75] on the PHOENIX14T dataset (Sec. B.5).
@@ B.3 Missing cue scenario
| Model | Dataset | Per-instance | Per-class | ||
@@
B.4 ISLR performance on How2Sign
@@
| Model | B1 | B2 | B3 | B4 | Rouge |
@@
B.5 Reproducing GFSLT and Sign2GPT on PHOENIX14T

### v2

@@ 2 Related Work
Continuous Sign Language Recognition (CSLR) typically involves recognising gloss sequences – spoken language words corresponding to semantic labels for individual signs – from a continuous signing video clip. It is crucial to distinguish CSLR from translation due to the significant grammatical differences between signed and spoken languages [69]. Benchmarks such as PHOENIX14 [41] and CSL-Daily [88], which have manually glossed training data, are widely used for CSLR. Due to a lack of temporal alignment between signs and video frames, many studies [21, 19, 76, 37, 11, 90, 32, 60, 53, 86] utilise the Connectionist Temporal Classification (CTC) loss [29]. More recently, [62] demonstrates the advantages of using video-to-text retrieval on the challenging BOBSL dataset [3].
@@
Recent works also incorporate large-scale language foundation models. For example, [68] inputs pseudo-glosses directly into ChatGPT. SignLLM [28] transforms sign language videos into discrete tokens, which are then fed into a frozen language model (LLaMA-7B-16bit [72]). [73, 64, 63, 38, 85] all fine-tune a pre-trained T5 model for SLT. [73] feeds in 3D landmark embeddings, while [64] and [63] use visual features from a BEVT-pre-trained [75] ISLR model and a MAE-pretained video encoder, respectively. Sign2GPT [78] leverages large-scale pre-trained vision (DINOv2 [57]) and language (XGLM [49]) models, incorporating adapters (LoRA [34]) for transfer to sign language. Similarly to these works, we make use of a pre-trained LLM (Llama3 [25]) and fine-tune it for SLT.
@@ 3.2 Model inputs
Visual features. Following [62, 59], we employ a Video-Swin model [50] pre-trained for ISLR (classification of isolated signs) to obtain our visual features. Specifically, we utilise the recent and strong model provided by [62], which is trained on the BOBSL videos [3] for a vocabulary of 8,697 signs using the automatic sign spottings from [54]. This ISLR model processes short video clips of 16 frames (i.e. less than 1 second in 25 fps videos) to produce a single 768-dimensional embedding vector. Specifically, this vector comes from spatio-temporal averaging of the penultimate layer output of Video-Swin. To capture fine-grained temporal details, we feed 16 consecutive frames into the sign video encoder and apply a sliding window with a stride of to obtain features , where represents the number of visual features per sentence. For example, when =2, we have on average 56 features (i.e. 4.5 seconds) in BOBSL sentences. Note that for experiments conducted on the How2Sign dataset [23], we further fine-tune the Video-Swin model using the spotting annotations provided by [24] for a vocabulary of 1,887 signs (see Sec. A.5 for further details).
Pseudo-glosses. To represent sign sequences in text form, we apply the ISLR model mentioned above in a sliding window manner to record the classification predictions, and obtain pseudo-glosses , representing a sequence of words (or phrases) from the vocabulary of the classifier (e.g. 8k signs). Note that these sign category predictions are noisy, often including more labels than the number of signs occurring in the sentence video, and many false positives, which we wish to suppress via our LLM tuning. For example, Fig. 2 shows an example of a homonym confusion between the manually similar signs ‘study’ and ‘worship’. Our pseudo-glosses are similar to [62], except we only apply repetition grouping, but do not filter out low-confidence annotations with a threshold – this allows the LLM to learn which ones are relevant. In contrast to [62] that uses pseudo-glosses for supervision, we simply employ them as additional inputs to our SLT model. Typically, there are around 22 glosses per BOBSL sentence (which is less than the number of visual features – 56 features on average).
@@ 3.3 Tuning the LLM with multifarious cues
Mapping network for visual features. All our inputs are already in text form except the visual features, which need a projection to map them into the text space, so that they can be fed into the pre-trained LLM. To this end, we train a simple mapping network, a 2-layer MLP with GELU [33] activation in between, projecting the visual features (of dimensionality 768) to the dimensionality of the LLM input embeddings (i.e. 4,096). We note that we add 1D temporal convolution layer to the mapping network when experimenting with the How2Sign dataset [23] (see Sec. A.5).
@@ 4 Experiments
In this section, we first present the datasets and a suite of evaluation protocols used in our experiments (Sec. 4.1), as well as baseline descriptions (Sec. 4.2). Next, we ablate various components of our framework (Sec. 4.3). We then show our improved translation performance compared to the state of the art on two challenging open-vocabulary benchmarks (Sec. 4.4). Finally, we illustrate qualitative results and discuss limitations (Sec. 4.5). Further experiments can be found in Appendix B.
@@ 4.2 Baselines
Oracle: Sincan [67] (Vid+Prev). This oracle baseline is a multi-modal variant of Sincan [67], where the previous ground truth sentence is also fed as additional context to the decoder at both training and inference times.
@@ 4.3 Ablation study
| Ours (Vid) | 2.6 | 37.8 | 15.6 | 37.5 | 13.6 | 0.95 |
| Extra cues | ||||||
| Ours (Vid+Prev) | 2.8 | 38.9 | 16.0 | 37.8 | 13.8 | 1.02 |
| Ours (Vid+Prev+PG) | 2.9 | 39.2 | 16.1 | 38.9 | 14.1 | 1.15 |
@@ 4.4 Comparison to the state of the art
BOBSL. We evaluate our model on the BOBSL test set, Sent-Test, using our full suite of evaluation metrics. As shown in Tab. 3, our approach achieves a significant improvement across all metrics compared to previous works. Notably, even with only video input, our model surpasses state-of-the-art methods such as GFSLT [87] and Sign2GPT [78]. Moreover, incorporating additional cues leads to a steady performance gain, with the full set of cues yielding a considerable boost (40.3 vs 37.8 B-RT). This highlights both the effectiveness of our method, which leverages context and the increased challenge posed by the BOBSL dataset (as opposed to PHOENIX14T where [87, 78] were originally evaluated).
In the Oracle setup (the bottom block of Tab. 3), we compare to the setting of [67], where models have access to ground truth previous sentence and spottings extracted from the ground truth current sentence. When using the ground truth previous sentence at inference, our model outperforms [67] by a large margin (40.5 vs 35.8 B-RT). When using both the ground truth previous sentence and spottings (which are obtained with access to current ground truth), we further increase the margin, substantially outperforming their method (45.9 vs 37.0 B-RT). Additionally, when integrating background descriptions (last row), we observe a further performance gain.
@@ 5 Conclusion
In this work, we show that leveraging contextual information significantly enhances SLT performance in an open-vocabulary setting. Specifically, our framework utilises background descriptions from a captioning model and predictions of previous sentences, combined with pseudo-glosses and visual features. Through an extensive ablation study, we analyse the individual impact of each cue on sign language translation, and benchmark our method against previous state-of-the-art approaches on the BOBSL dataset to demonstrate its effectiveness. Despite the improvements of our approach, several challenges remain. Background descriptions may not always complement the translation, potentially introducing noise. Relying on previous sentence predictions can lead to error accumulation, where early mistakes propagate through subsequent sentences. Additionally, pseudo-glosses are prone to false positives due to homonyms, resulting in incorrect word choices. Addressing these limitations would potentially improve the robustness and real-world applicability of sign language translation systems.
Acknowledgments.
The images in this paper are used with kind permission of the BBC. This work was granted access to the HPC resources of IDRIS under the allocation 2024-AD011013395 made by GENCI. The authors would like to acknowledge the ANR project CorVis ANR-21-CE23-0003-01, a Google Research Scholar Award, and a Royal Society Research Professorship RPR1191132. The authors also thank Charles Raude, Prajwal KR, Prasanna Sridhar, Joon Son Chung, Makarand Tapaswi, Syrine Kalleli, and David Picard for their feedback, and further thank Ryan Wong for guiding the Sign2GPT [78] implementation and Ozge Mercanoglu Sincan for providing the outputs for the Sincan [67] model on the BOBSL test set.
@@ Appendix B Additional Experiments
We examine performance variations when using different LLM decoders (Sec. B.1), evaluate all possible cue combinations (Sec. B.2), investigate scenarios with missing cues (Sec. B.3), with multiple number of previous sentences (Sec. B.4), and with various background frame sampling rates (Sec. B.5). We also showcase the performance of our ISLR backbone on the HowSign dataset (Sec. B.6), and report the reproduction results of GFSLT [87] and Sign2GPT [78] on the PHOENIX14T dataset (Sec. B.7). Finally, we demonstrate the applicability of our method on PHOENIX14T (Sec. B.8).
@@ B.3 Missing cue scenario
| #prev. | B4 | B-RT | R-L | CIDEr | IoU | LLM |
| 1 | 3.3 | 40.3 | 16.9 | 41.9 | 14.8 | 1.20 |
| 2 | 3.3 | 40.6 | 17.0 | 42.8 | 14.9 | 1.21 |
| 3 | 3.4 | 41.0 | 17.2 | 43.9 | 15.1 | 1.24 |
B.4 Number of previous sentences
Tab. A.7 reports the results of our best model (Vid+PG+Prev +BG) on Sent-Test, using 2–3 previous sentences as context during both training and inference. As shown in the table, providing a longer previous context results in only a marginal improvement (+0.1 BLEU-4 and +0.3 ROUGE). While additional context slightly improves B-RT (41.0 vs. 40.3), it comes at the cost of increased computational overhead during inference. To balance performance and efficiency, we use only a single previous sentence for inference.
| Sampling rate (sec) | B4 | B-RT | R-L | CIDEr | IoU | LLM |
| 1 | 3.3 | 40.3 | 16.9 | 41.9 | 14.8 | 1.20 |
| 2 | 3.3 | 40.2 | 17.0 | 41.3 | 14.8 | 1.20 |
| 3 | 3.1 | 40.0 | 16.9 | 40.9 | 14.5 | 1.19 |
B.5 Background frame sampling rate
Tab. A.8 reports the results of our best model (Vid+PG+Prev +BG) on Sent-Test, when varying the sampling rate of the background frames during both training and inference. In the rest of the experiments, we sample a background caption every 1 second. In Tab. A.8, we experiment with reducing this rate by sampling every 2 or 3 seconds. Since we remove repeated words in background captions, we do not experiment with sampling more than 1 frame per second (adjacent frames often depict nearly identical scenes). We obtain (40.3, 40.2, 40.0) BLEURT when sampling every (1, 2, 3) seconds, indicating little effect on performance at lower sampling rates. Thus, sampling every second avoids missing scene transitions while remaining computationally feasible.
| Model | Training | Per-instance | Per-class | ||
@@ B.5 Background frame sampling rate
B.6 ISLR performance on How2Sign
@@
B.7 Reproducing GFSLT and Sign2GPT on PHOENIX14T
| Model | B1 | B2 | B3 | B4 | R-L |
@@ B.7 Reproducing GFSLT and Sign2GPT on PHOENIX14T
| Model | B4 | B-RT | R-L | CIDEr | IoU |
| GFSLT [87] | 21.44 | - | 42.49 | - | - |
| Sign2GPT [78] | 19.42 | - | 45.23 | - | - |
| Sign2GPT (w/PGP) [78] | 22.52 | - | 48.90 | - | - |
| Ours (Vid) | 20.58 | 52.25 | 41.20 | 190.40 | 34.05 |
| Ours (Vid+PG [1]) | 23.80 | 52.80 | 46.11 | 227.05 | 38.49 |
@@
B.8 Evaluation of our model on PHOENIX14T
To evaluate the generalisability of our model, we conduct experiments on the PHOENIX14T dataset [41, 12]. We extract video features using a Video-Swin model trained with pseudo-glosses (PG) obtained from SlowFastSign [1]. As shown in Tab. A.11, when fine-tuning an LLM-based model using only video features, we achieve a BLEU-4 score of 20.58 and a ROUGE-L score of 41.20. By incorporating PG as an additional textual cue, performance improves to a BLEU-4 score of 23.80 and a ROUGE-L score of 46.11. Compared to Sign2GPT [78], which reports a BLEU-4 score of 22.52 and a ROUGE-L score of 48.90, our model demonstrates comparable performance while highlighting the effectiveness of leveraging PG as an additional cue.
Note that no other contextual information (i.e. previous sentence, background) is available for this evaluation. Although PHOENIX14T consists of TV weather broadcasts, the dataset is segmented at the sentence level and lacks context from previous sentences or background information. Moreover, the dataset includes manually annotated glosses and has a restricted vocabulary, leading to performance saturation. This setting therefore does not reflect open-vocabulary tasks, which present greater challenges in real-world scenarios.